OpenStudy (1230kent):

In Jim's class, 5 times as many students take the bus as those who walk. A total of 24 students walk or take the bus. How many more students take the bus than walk to school? 5, 6, 16, 20 This for my 4th grader so please explain so I can show her.

OpenStudy (1230kent):

I think the answer is 20 because 5*4=20 with 4 left over

OpenStudy (anonymous):

Let B is number of student take Bus W ............ Walk then B = 5W -> 5W + W = 24 Students

OpenStudy (anonymous):

Yes, good guess ;)

OpenStudy (1230kent):

Thank you so much for the explaination of the question

OpenStudy (anonymous):

So you can finish the rest?

OpenStudy (1230kent):

no, 5W+W=24 6w=24 w=4 4 walkers and 20 bus riders

OpenStudy (anonymous):

You mean "Yes", right!

OpenStudy (1230kent):

I did that right

OpenStudy (anonymous):

If it's for 4 grader, look like just use trial and error method: 5n + n = 24 Let n = 1, 2, ..6 until you gain the combination = 24!

OpenStudy (1230kent):

Can you help me with one more?

OpenStudy (anonymous):

Just open the new post for each question :)

OpenStudy (1230kent):

Kirin has 28 books. This is 7 times as many books as Gail has. Kirin made a model to compare the numbers of books they have. 28 n n n n n n n n which equation represents how to find the vakue of n 28=n+7 28=7*n 28+7=n 28*7=n

OpenStudy (anonymous):

Okay, the question asks how many MORE students take the bus as opposed to walking. So, this equation is apt (n being the number of students who walk, 5n representing the kids on a bus): 5n + n = 24. So 6n = 24. Then that means that 6(4) = 24. n = 4. This means four kids walk to school, while 20 take the bus. And how many more students bus as opposed to walk is solved by the number sentence: 20 (which is 5n; bus kids) - 4 (n; walkers) = 16 more kids ride bus than walk. The answer is 16. You can explain this to a 4th grader this way by using a model for an unknown number: bus kids n + n + n + n + n (5 times as many walkers or "n") walkers n So, six "n"s, = 24 total children who walk or bus to school. 6 x n = 24. Using multiplication facts to solve this equation you can see that 6 x 4 = 24. So, the number of kids who walk is 4, and the number of kids who bus (24 - 4 = 20). Then you have to find many more students bus than walk, and you solve that with simple subtraction: 20 - 4 = 16. I know I'm two years too late, but I had to post a correction to all the responses I just read. I really hope this is helpful to someone in the future.